OpenStudy (anonymous):

I have a few problems that I need desperate help on! Can anyone help? 1) f(x)=10/7x; g(x)=7/x What is the domain of (2) What is the domain of (7) 2.) Ln,g(x)=e^2x and h(x)=x^5 A) Fog (x) Domain B) Gof (x) Domain C) Foh (x) Domain

OpenStudy (anonymous):

f(x) , g(x) FoG = F(g(x)) Gof= g(f(x)) .. so in other word your still finding a domani the normal way except were you would normal have x or y plug the designated function into the variable spot first then solve for domain.

OpenStudy (anonymous):

so for example y=f(x)= 2x+4 and g(x)=3x fog --> f(g(x))= 2(g(x))+4 fog=6x+4

OpenStudy (anonymous):

can you show me how to do mine?

OpenStudy (anonymous):

do you understand what i showed you?

OpenStudy (anonymous):

I am so frustrated with this I have no clue. I have a huge test is due tomorrow.. beyond frustrated

OpenStudy (anonymous):

you know what domain and range are?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

f(x)=10/7x g(x)=7/x so FoG is f(g(x)) do you understand that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

gof be... g(f(x))

OpenStudy (anonymous):

right so what is Fog how would you do it?

OpenStudy (anonymous):

f(h(x))

OpenStudy (anonymous):

ok well fog= \[10/7(g(x))\]

OpenStudy (anonymous):

g(x)=7/x

OpenStudy (anonymous):

=10x/49

OpenStudy (anonymous):

when they domain of (20 what are they referring too

OpenStudy (anonymous):

*(2)

OpenStudy (anonymous):

Idk